What mass of sucrose (C 12 H 22 O 11 ) is dissolved in 50.0 mL of a 0.400 mol/L solution of
sucrose in water?
1
Expert's answer
2020-05-10T09:19:24-0400

n(C12H22O11) = V*C = 0.050*0.400 = 0.02 mol

m(C12H22O11) = m*M = 0.02*342 = 6.84 g